# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

### v1

@@ A.1 Useful algebriac facts
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})=\bm{I}-\eta(\bm{A}+\bm{A}^{\top})+\eta^{2}\bm{A}^{\top}\bm{A},$
where $\bm{A}^{\top}\bm{A}$ can be bounded above by
$\bm{A}^{\top}\bm{A}$ $\preceq\bm{A}^{\top}\bm{A}+\bm{A}\bm{A}^{\top}\preceq\frac{1}{2}(\bm{A}+\bm{A}^{\top})^{2}\preceq(1+\gamma)\lambda_{\Sigma}(\bm{A}+\bm{A}^{\top}),$
where we applied Equation (30) in the last inequality. Hence,
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-(\eta-(1+\gamma)\lambda_{\Sigma}\eta^{2})(\bm{A}+\bm{A}^{\top}).$
For any $\eta\in(0,\frac{1+\gamma}{2\lambda_{\Sigma}})$, this upper bound is controlled by
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-\frac{\eta}{2}(\bm{A}+\bm{A}^{\top})\preceq(1-(1-\gamma)\lambda_{0})\bm{I}.$
@@
$\|\bm{I}-\eta\bm{A}\|\leq\sqrt{1-(1-\gamma)\lambda_{0}}\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta.$
@@
For every positive interger $t$, we begin by identifying the index $i$ that maximizes
@@
$\sum_{j=t}^{\infty}(j+1)^{-\alpha}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$ $=\frac{1}{\beta}\sum_{j=t}^{\infty}\left(\prod_{k=t+1}^{j}-\prod_{k=t+1}^{j+1}\right)(1-\beta k^{-\alpha})$
$=\frac{1}{\beta}\left(1-\prod_{k=t+1}^{\infty}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$
@@
Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta_{0}\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$. Recall that $\bm{Q}_{t}$ is defined as in (19). Then for any positive integer $\ell<T$, the following relation hold:
@@
By definition, the sum of $Q_{t}$ for $1\leq t\leq T$ can be expressed as
$\sum_{t=1}^{T}\bm{Q}_{t}=\sum_{t=1}^{T}\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$
By switching the order of summation we obtain
@@
$=\sum_{j=1}^{T}\sum_{t=1}^{j}\bm{A}^{-1}\left(\prod_{k=t+1}^{j}-\prod_{k=t}^{j}\right)(\bm{I}-\eta_{k}\bm{A})$
@@
Under the condition 12, for every positive definite matrix $\bm{Y}\in\mathbb{R}^{d\times d}$, it can be guarnateed that
@@
$T^{-\alpha}\sum_{t=1}^{T}\left(\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right)=\bm{X}+\bm{O}(T^{\alpha-1}),$ (51)
where $\bm{X}$ is the solution to the Lyapunov equation
$\eta_{0}(\bm{A}\bm{X}+\bm{XA}^{\top})=\bm{Y}.$ (52)
@@
Proof: For simplicity, denote the left-hand-side of (51) as $\bm{X}_{T}$. It can be easily verified that the sequence $\{\bm{X}_{T}\}_{T\geq 1}$ is featured by the iterative relation
$\bm{X}_{T}$ $=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{X}_{T-1}(\bm{I}-\eta_{T}\bm{A}^{\top})$
$+T^{-\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{Y}(\bm{I}-\eta_{T}\bm{A}^{\top}).$
Direct calculation yields
$\bm{X}_{T}-\bm{X}$ $=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})(\bm{X}_{T-1}-\bm{X})(\bm{I}-\eta_{T}\bm{A}^{\top})$
$+\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{X}(\bm{I}-\eta_{T}\bm{A}^{\top})-\bm{X}$
$+T^{-\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{Y}(\bm{I}-\eta_{T}\bm{A}^{\top})$
$=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})(\bm{X}_{T-1}-\bm{X})(\bm{I}-\eta_{T}\bm{A}^{\top})$
$+T^{-\alpha}[\eta_{0}(\bm{AX}+\bm{XA}^{\top})-\bm{Y}]+\frac{\bm{X}}{T}+O(T^{-2\alpha}).$
Notice that the first term on the last line is $\bm{0}$ according to the definition of $\bm{X}$; therefore by iteration, the difference between $\bm{X}_{T}$ and $\bm{X}$ can be represented by
$\bm{X}_{T}-\bm{X}$ $=T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})$
$-T^{-\alpha}\prod_{k=1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+O(T^{-\alpha});$
By letting $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$, LemmaA.3 inplies the following bound on the norm of the first term on the right-hand-side:
$\left\|T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$
$<T^{\alpha-1}\sum_{i=1}^{T}i^{-\alpha}\prod_{k=1}^{T}(1-\beta k^{-\alpha})\|\bm{X}\|$
$=T^{\alpha-1}\cdot\frac{1}{\beta}\sum_{i=1}^{T}\left(\prod_{k=i}^{T}-\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})\right)\|\bm{X}\|<\frac{\|\bm{X}\|}{\beta}T^{\alpha-1}.$
The two terms on the second line are clearly both $\bm{o}(T^{\alpha-1})$; hence,
$\|\bm{X}_{T}-\bm{X}\|=\bm{O}(T^{\alpha-1}).$
@@
Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$. For $\bm{Q}_{t}$ defined as in (19), it can be guaranteed that
@@
Proof: According to Lemma A.3, since $0<\eta_{t}\leq\eta_{0}\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$ for all $t$, the norm of $\bm{Q}_{t}$ is bounded by
@@
$\|\bm{AQ}_{t}-\bm{I}\|^{2}\leq\|\bm{AS}_{t}\|^{2}+\left\|\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\right\|^{2}$ $\leq\widetilde{C}^{\prime}t^{2\alpha-2}+\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
@@
$=\widetilde{C}^{\prime}t^{2\alpha-2}+T^{\alpha}\frac{1}{\beta}\left(\prod_{k=t}^{T}-\prod_{k=t-1}^{T}\right)(1-\beta k^{-\alpha});$
The bound (55) follows by taking an average from $t=1$ to $t=T$ and applying the telescoping method. ∎
@@
then as $T\to\infty$, the difference between $\overline{\bm{\Lambda}_{T}}$ and $\bm{\Lambda}^{\star}$ is featured by
$T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})-\bm{O}(T^{2\alpha-2})\preceq\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}\preceq T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+\bm{O}(T^{2\alpha-2})$
Proof: By definition, the difference bewteen $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ can be expressed as
@@
where we applied Equation (49a) and the interchangability of the production. Similarly, the second term can be expressed by
@@
Lemma A.3 directly implies that the norm of these two terms are bounded by
$\left\|\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}\right\|$ $\leq\frac{1}{T}\|\bm{\Lambda}^{\star}\|\sum_{j=1}^{T}\prod_{k=1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{0}k^{-\alpha}\right).$
@@
a constant independent of $T$. Therefore, $I_{1}$ is bounded above by $\bm{O}(T^{-1})$.

### v3

@@ A.1 Useful algebraic facts
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})=\bm{I}-\eta(\bm{A}+\bm{A}^{\top})+\eta^{2}\bm{A}^{\top}\bm{A}.$
By invoking (29), we obtain
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-(\eta-\lambda_{\Sigma}\eta^{2})(\bm{A}+\bm{A}^{\top}).$
For any $\eta\in(0,\frac{1}{2\lambda_{\Sigma}})$, this upper bound is controlled by
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-\frac{\eta}{2}(\bm{A}+\bm{A}^{\top})\preceq(1-(1-\gamma)\lambda_{0}\eta)\bm{I}.$
@@
$\|\bm{I}-\eta\bm{A}\|\leq\sqrt{1-(1-\gamma)\lambda_{0}\eta}\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta.$
Proof of Equation (31).
For any unit vector $\bm{x}\in\mathbb{R}^{d}$, we observe
$\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}$ $=\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s)]\bm{\Sigma}^{-\frac{1}{2}}\bm{x}-\gamma\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s^{\prime})]\bm{\Sigma}^{-\frac{1}{2}}\bm{x}$
$=\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\Sigma}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}-\gamma\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s^{\prime})]\bm{\Sigma}^{-\frac{1}{2}}\bm{x}$
$=1-\gamma\mathbb{E}[(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s))(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s^{\prime}))]$
$\leq 1-\gamma\sqrt{\mathbb{E}[(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s))^{2}]\mathbb{E}[(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s^{\prime}))^{2}]}$
$=1-\gamma\sqrt{(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s)]\bm{\Sigma}^{-\frac{1}{2}}\bm{x})\cdot(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s^{\prime})\bm{\phi}^{\top}(s^{\prime})]\bm{\Sigma}^{-\frac{1}{2}}\bm{x})}$
$=1-\gamma\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\Sigma}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}=1-\gamma,$
where the fourth line is an application of the Cauchy-Schwartz inequality. Consequently, for any unit vector $\bm{x}\in\mathbb{R}^{d}$, it can be guaranteed that
$\|\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}\|_{2}$ $=\|\bm{x}\|_{2}\|\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}\|_{2}$
$\geq\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}\geq 1-\gamma,$
where the inequality, again, is an application of the Cauchy-Schwartz inequality. This essentially means that $\|\bm{\Sigma}^{\frac{1}{2}}\bm{A}^{-1}\bm{\Sigma}^{\frac{1}{2}}\|\leq\frac{1}{1-\gamma}$, and therefore
$\|\bm{A}^{-1}\|\leq\frac{1}{1-\gamma}\frac{1}{\lambda_{\min}(\bm{\Sigma})}=\frac{1}{\lambda_{0}(1-\gamma)}.$
Lemma A.4 implies the following property regarding a Lyapunov equation.
Lemma A.5.
For any matrix $\bm{E}\succeq\bm{0}$, there exists a unique positive definite matrix $\bm{X}\succeq\bm{0}$, such that
$\bm{AX}+\bm{XA}^{\top}=\bm{E}.$
Furthermore, it can be guaranteed that
$\|\bm{X}\|\leq\frac{1}{2(1-\gamma)\lambda_{0}}\|\bm{E}\|,\quad\text{and}\quad\mathsf{Tr}(\bm{X})\leq\frac{1}{2(1-\gamma)\lambda_{0}}\mathsf{Tr}(\bm{E}).$
Proof: The existence of $\bm{X}$ follows from the basic property of Lyapunov equation and (27); we focus on bounding the norm and trace of $\bm{X}$.
Bounding the operator norm of $\bm{X}$.
Let $\lambda=\lambda_{\max}(\bm{X})$ and $\bm{x}$ denote a corresponding eigen vector with $\|\bm{x}\|_{2}=1$. By definition, $\bm{Xx}=\lambda\bm{x}$, and therefore
$\|\bm{E}\|\geq\bm{x}^{\top}\bm{Ex}=\bm{x}^{\top}(\bm{AX}+\bm{XA}^{\top})\bm{x}=\bm{x}^{\top}(\bm{A}+\bm{A}^{\top})\bm{x}\geq[2(1-\gamma)\lambda_{0}]\cdot\lambda,$
where the last inequality follows from (26). The bound on $\|\bm{X}\|$ follows naturally.
Bounding the norm of $\bm{X}$
. Since $\bm{X}\succeq\bm{0}$, direct calculation yields
$\mathsf{Tr}(\bm{E})$ $=\mathsf{Tr}(\bm{AX}+\bm{XA}^{\top})$
$=\mathsf{Tr}(\bm{X}^{\frac{1}{2}}\bm{AX}^{\frac{1}{2}}+\bm{X}^{\frac{1}{2}}\bm{A}^{\top}\bm{X}^{\frac{1}{2}})$
$=\mathsf{Tr}(\bm{X}^{\frac{1}{2}}(\bm{A}+\bm{A}^{\top})\bm{X}^{\frac{1}{2}})$
$\geq\mathsf{Tr}(\bm{X}^{\frac{1}{2}}\cdot 2(1-\gamma)\lambda_{0}\cdot\bm{X}^{\frac{1}{2}})=2(1-\gamma)\lambda_{0}\mathsf{Tr}(\bm{X}),$
where we invoked the cyclability of trace and (26) through the process. ∎
@@
For every positive integer $t$, we begin by identifying the index $i$ that maximizes
@@
$\sum_{j=t}^{\infty}(j+1)^{-\alpha}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$
$=\sum_{j=t}^{\infty}\frac{1}{\beta}\left[1-(1-\beta(j+1)^{-\alpha})\right]\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$
$=\frac{1}{\beta}\sum_{j=t}^{\infty}\left(\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})-\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)$
$=\frac{1}{\beta}\left(\sum_{j=t}^{\infty}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)$
$=\frac{1}{\beta}\left[\left(\prod_{k=t+1}^{t}(1-\beta k^{-\alpha})+\sum_{j=t+1}^{\infty}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})\right)-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right]$
=$\overset{(i)}{=}\frac{1}{\beta}\left(1+\sum_{i=t}^{\infty}\prod_{k=t+1}^{i+1}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$ (45)
Notice that in (i) we invoked the convention
$\prod_{k=t+1}^{t}(1-\beta k^{-\alpha})=1,$
and a change of variable $i=j-1$.
@@
Next, we introduce additional quantities
where $\bm{\Gamma}$ is defined in (12) and, for each $t$, we set
$\bm{Q}_{t}=\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$ (50)
The matrix $\overline{\bm{\Lambda}}_{T}$ plays the role of a non-asymptotic covariance matrix for $\overline{\bm{\theta}}_{T}$. The following lemmas feature the properties of $\bm{Q}_{t}$ and $\overline{\bm{\Lambda}}_{T}$.
@@
Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta_{0}\leq\frac{1}{2\lambda_{\Sigma}}$. Recall that $\bm{Q}_{t}$ is defined as in (50). Then for any positive integer $\ell<T$, the following relation hold:
@@
By definition, the sum of $\bm{Q}_{t}$ for $1\leq t\leq T$ can be expressed as
$\sum_{t=1}^{T}\bm{Q}_{t}$ $=\sum_{t=1}^{T}\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$
$=\sum_{j=1}^{T}\sum_{t=1}^{j}\eta_{t}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}),$
where the second line follows by switching the order of summation. Here, the summand can be represented as
$\eta_{t}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$ $=\bm{A}^{-1}[\bm{I}-(\bm{I}-\eta_{t}\bm{A})]\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$
$=\bm{A}^{-1}\left[\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})-\prod_{k=t}^{j}(\bm{I}-\eta_{k}\bm{A})\right].$ (52)
Notice that summing the difference in the bracket from $t=1$ to $t=j$ would cancel all the terms in the middle and only the first and last terms would remain. Therefore,
@@
$=\sum_{j=1}^{T}\sum_{t=1}^{j}\bm{A}^{-1}\left[\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})-\prod_{k=t}^{j}(\bm{I}-\eta_{k}\bm{A})\right]$
@@
When $\eta_{0}<\frac{1}{2\lambda_{\Sigma}}$, for every positive definite matrix $\bm{Y}\in\mathbb{R}^{d\times d}$, it can be guaranteed that
$T^{-\alpha}\sum_{t=1}^{T}\left(\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right)-\bm{X}\leq O(T^{\alpha-1})\bm{Y}$ (54)
where $\bm{X}\succeq\bm{0}$ is the solution to the Lyapunov equation
$\eta_{0}(\bm{A}\bm{X}+\bm{XA}^{\top})=\bm{Y},$ (55)
and the big $O$ notation for matrix sequences is defined as in (22).
Proof: For simplicity, we define a sequence of matrices $\{\bm{X}_{i}\}_{i=0}^{T}$ as
@@
$\bm{X}_{i}=\begin{cases}&i^{-\alpha}\sum_{t=1}^{i}\left(\prod_{k=t}^{i}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=t}^{i}(\bm{I}-\eta_{k}\bm{A}^{\top})\right),\quad\text{for}\quad 1\leq i\leq T;\\ &\bm{0},\quad\text{for}\quad i=0.\end{cases}$
@@
in this way, the left-hand-side of (54) can be represented by $\bm{X}_{T}-\bm{X}$. It can be easily verified that the sequence $\{\bm{X}_{i}\}_{i=1}^{T}$ is featured by the iterative relation
$i^{\alpha}\bm{X}_{i}$ $=(i-1)^{\alpha}(\bm{I}-\eta_{i}\bm{A})\bm{X}_{i-1}(\bm{I}-\eta_{i}\bm{A}^{\top})+(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top}),\quad\forall 1\leq i\leq T.$
Direct calculation yields that for every $1\leq i\leq T$,
$i^{\alpha}(\bm{X}_{i}-\bm{X})$ $=(i-1)^{\alpha}(\bm{I}-\eta_{i}\bm{A})\bm{X}_{i-1}(\bm{I}-\eta_{i}\bm{A}^{\top})+(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top})-i^{\alpha}\bm{X}$
$=(i-1)^{\alpha}(\bm{I}-\eta_{i}\bm{A})(\bm{X}_{i-1}-\bm{X})(\bm{I}-\eta_{i}\bm{A}^{\top})$
$+(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top})-[i^{\alpha}-(i-1)^{\alpha}]\bm{X}-(i-1)^{\alpha}[\eta_{i}(\bm{AX}+\bm{XA}^{\top})+\eta_{i}^{2}\bm{AXA}^{\top}].$
We now feature the terms on the last line. Firstly, we notice that (55) guarantees
$\eta_{i}(\bm{AX}+\bm{XA}^{\top})=\eta_{0}i^{-\alpha}(\bm{AX}+\bm{XA}^{\top})=i^{-\alpha}\bm{Y};$
therefore, a simple reorganization yields
$(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top})-[i^{\alpha}-(i-1)^{\alpha}]\bm{X}-(i-1)^{\alpha}[\eta_{i}(\bm{AX}+\bm{XA}^{\top})+\eta_{i}^{2}\bm{AXA}^{\top}]$
$=\left[1-\left(\frac{i-1}{i}\right)^{\alpha}\right]\bm{Y}-[i^{\alpha}-(i-1)^{\alpha}]\bm{X}-\eta_{i}(\bm{AY}+\bm{YA}^{\top})+\eta_{i}^{2}\bm{A}(\bm{Y}-(i-1)^{\alpha}\bm{X})\bm{A}^{\top}.$
For these terms, we observe
$\left[1-\left(\frac{i-1}{i}\right)^{\alpha}\right]\bm{Y}\preceq\left[1-\left(\frac{i-1}{i}\right)\right]\bm{Y}=\frac{1}{i}\bm{Y};$
$[i^{\alpha}-(i-1)^{\alpha}]\bm{X}=i^{\alpha}\left[1-\left(\frac{i-1}{i}\right)^{\alpha}\right]\bm{X}\preceq i^{\alpha}\left[1-\left(\frac{i-1}{i}\right)\right]\bm{X}=i^{\alpha-1}\bm{X}=O(i^{\alpha-1})\bm{Y};$
$\eta_{i}(\bm{AY}+\bm{YA}^{\top})=O(i^{-\alpha})\bm{Y};$
$\eta_{i}^{2}\bm{A}(\bm{Y}-(i-1)^{\alpha}\bm{X})\bm{A}^{\top}=O(i^{-\alpha})\bm{Y}.$
Hence, the leading term decreases by the rate of $O(i^{\alpha-1})\bm{Y}$. As a result,
$i^{\alpha}(\bm{X}_{i}-\bm{X})=(\bm{I}-\eta_{i}\bm{A})[(i-1)^{\alpha}(\bm{X}_{i-1}-\bm{X})](\bm{I}-\eta_{i}\bm{A}^{\top})+O(i^{\alpha-1})\bm{Y}$
For simplicity, we denote $\bm{Z}_{i}=i^{\alpha}(\bm{X}_{i}-\bm{X})$. The relation above can then be represented as
$\bm{Z}_{i}=(\bm{I}-\eta_{i}\bm{A})\bm{Z}_{i-1}(\bm{I}-\eta_{i}\bm{A}^{\top})+O(i^{\alpha-1})\bm{Y}.$
It is easy to verify by iteration that, there exists a quantity $\widetilde{C}$ independent of $T$, such that
$\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})$
$\preceq\bm{Z}_{T}\preceq\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top});$
In other words,
$T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})$
$\preceq\bm{X}_{T}-\bm{X}$
$\preceq T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top}).$
By letting $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$, Lemma A.4 implies
$\left\|T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$
$\leq T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}\|\bm{I}-\eta_{k}\bm{A}\|\prod_{k=i+1}^{T}\|\bm{I}-\eta_{k}\bm{A}^{\top}\|$
$\leq T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})^{2}$
$<T^{\alpha-1}\sum_{i=1}^{T}i^{-\alpha}\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})$
$=T^{\alpha-1}\cdot\frac{1}{\beta}\sum_{i=1}^{T}\left(\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})-\prod_{k=i}^{T}(1-\beta k^{-\alpha})\right)$
$=\frac{1}{\beta}T^{\alpha-1}\left(\prod_{k=T+1}^{T}(1-\beta k^{-\alpha})-\prod_{k=1}^{T}(1-\beta k^{-\alpha})\right)<\frac{1}{\beta}T^{\alpha-1};$
notice that in the last two lines we invoked a similar approach to the telescoping method (A.1). Combining this with (25), we obtain
$T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})=O(T^{\alpha-1})\bm{Y}.$
This completes the proof. ∎
@@
Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta\leq\frac{1}{2\lambda_{\Sigma}}$. For $\bm{Q}_{t}$ defined as in (50), it can be guaranteed that
@@
Proof: According to Lemma A.4, since $0<\eta_{t}\leq\eta_{0}\leq\frac{1}{2\lambda_{\Sigma}}$ for all $t$, the norm of $\bm{Q}_{t}$ is bounded by
@@
$\|\bm{AQ}_{t}-\bm{I}\|^{2}$ $\leq\|\bm{AS}_{t}\|^{2}+\left\|\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\right\|^{2}$
$\leq\widetilde{C}^{\prime}t^{2\alpha-2}+\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
@@
$=\widetilde{C}^{\prime}t^{2\alpha-2}+T^{\alpha}\cdot\frac{1}{\beta}\left(\prod_{k=t}^{T}(1-\beta k^{-\alpha})-\prod_{k=t-1}^{T}(1-\beta k^{-\alpha})\right);$
The bound (58) follows by taking an average from $t=1$ to $t=T$ and applying the telescoping method as in (A.1). ∎
@@
then as $T\to\infty$, the difference between $\overline{\bm{\Lambda}_{T}}$ and $\bm{\Lambda}^{\star}$ satisfies
$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}-T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})=O(T^{2\alpha-2})\bm{\Gamma}.$
Here, the $O$ notation for matrices is defined as in (22).
Proof: By definition, the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ can be expressed as
@@
where we applied Equation (51a) the interchangeability of products. Similarly, the second term can be expressed by
@@
Lemma A.4 directly implies that
$\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\leq\frac{1}{T}\prod_{k=1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{0}k^{-\alpha}\right).$
@@
a constant independent of $T$. Therefore, $I_{1}$ is featured by
$\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}+\frac{1}{T}\sum_{t=1}^{T}\bm{A}^{-1}\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}=O(T^{-1})\bm{\Lambda}^{\star}.$